truck travelling at 22.5 m/s decelerates at 2.27 m/s2



. (a) How much time does it take for the truck to stop? [9.91 ]s



(b) How far does it travel while stopping? [112m]



(c) How far does it travel during the third second after the brakes are applied? [

1
Expert's answer
2022-05-18T08:32:02-0400

(a)

"v=v_0+at"

"t=\frac{v-v_0}{a}=\frac{0-22.5}{-2.27}=9.91\:\rm s"

(b)

"d=\frac{v^2-v_0^2}{2a}=\frac{0^2-22.5^2}{2*(-2.27)}=112\:\rm m"

(c)

"\Delta d=d_3-d_2\\
=(22.5*3-2.27*3^2/2)\\
-(22.5*2-2.27*2^2/2)\\
=16.8\:\rm m"